An unknown element (X) has two stable isotopes. 18X with a mass of 17.55 amu: has a % abundance of 83.55% while the other isotope is (19X) whose mass is 18.67 amu has a % abundance of 16.45%. What is the average and relative atomic mass of the unknown element?
1
Expert's answer
2020-09-30T07:50:14-0400

Solution:

1) The average atomic mass of an element is the sum of the masses of its isotopes, each multiplied by its natural abundance.

Average atomic mass = f1M1 + f2M2 +… + fnMn, where f is the fraction representing the natural abundance of the isotope and M is the mass number (weight) of the isotope.

Thus:

Average atomic mass of X = (0.8355) × (17.55 amu) + (0.1645) × (18.67 amu) = 17.73 amu


2) The relative atomic mass is the average of the mass numbers of the different isotopes.

The relative atomic mass of X = (0.8355) × (18) + (0.1645) × (19) = 18.1645